# CVC: A Large-Scale Chinese Value Rule Corpus for Value Alignment of Large Language Models

### v4

@@ Abstract
Ensuring that Large Language Models (LLMs) align with mainstream human values and ethical norms is crucial for the safe and sustainable development of AI. Current value evaluation and alignment are constrained by Western cultural bias and incomplete domestic frameworks reliant on non-native rules; furthermore, the lack of scalable, rule-driven scenario generation methods makes evaluations costly and inadequate across diverse cultural contexts. To address these challenges, we propose a hierarchical value framework grounded in core Chinese values, encompassing three main dimensions, 12 core values, and 50 derived values. Based on this framework, we construct a large-scale Chinese Values Corpus (CVC) containing over 250,000 value rules enhanced and expanded through human annotation. Experimental results show that CVC-guided scenarios outperform direct generations ones in value boundaries and content diversity. In the evaluation across six sensitive themes (e.g., surrogacy, suicide), seven mainstream LLMs preferred CVC-generated options in over 70.5% of cases, while five Chinese human annotators showed an 87.5% alignment with CVC, confirming its universality, cultural relevance, and strong alignment with Chinese values. Additionally, we construct 400,000 rule-based moral dilemma scenarios that objectively capture nuanced distinctions in conflicting value prioritization across 17 LLMs. Our work establishes a culturally-adaptive benchmarking framework for comprehensive value evaluation and alignment, representing Chinese characteristics. All data are available at https://huggingface.co/datasets/Beijing-AISI/CVC, and the code is available at https://github.com/Beijing-AISI/CVC.
1 Introduction
In recent years, Large Language Models (LLMs) have been widely deployed across various domains, exerting growing influence on human decision-making and behaviortessler2024ai ; hu2025generative . However, their outputs still pose significant risks, including harmful biashu2025generative , hallucinationsun2024ai ; zhou2024larger , and factual inconsistencylin2022truthfulqa . These risks highlight the urgent need for effective evaluation frameworks that can assess and guide the ethical behavior of LLMs, ensuring that their responses remain aligned with prevailing societal values and moral expectations.(shen2023large, ; taddeo2018ai, ).
@@ 1 Introduction
To address these issues, in this paper, we propose a hierarchical value framework based on core Chinese values. With the assistance of LLMs and manual verification, we constructed a large-scale, carefully curated and standardized Chinese values corpus containing over 250,000 rules. The proposed CVC can effectively guide the generation of value assessment scenarios, demonstrating significant advantages in thematic relevance, value boundaries, content diversity, and semantic clarity. Furthermore, with high-quality human annotation ensuring alignment with Chinese cultural context, CVC provides a benchmark that both represents local values and connects with the universal ethical principles recognized by existing LLMs. Compared to other benchmarks, it provides more localized and precise standards for value alignment. Finally, CVC enables automated construction of complex, large-scale moral dilemma scenarios that systematically avoid biases while capturing the trade-offs between conflicting values.
@@
    The first large-scale, curated Chinese Values Corpus (CVC): Based on the core socialist values, we develop a localized value framework covering national, societal, and personal dimensions, with 12 core values and 50 derived values. Building upon this framework, we build the first large-scale Chinese values corpus, comprising over 250,000 high-quality, manually annotated and augmented normative rules.
@@
    Systematic validation of CVC’s generative advantages and value alignment: We validated the effectiveness of CVC across all 12 core value dimensions, where CVC-guided scenarios demonstrated clearer semantic boundaries and better category separation.Notably, CVC-guided scenarios in categories such as rule of law and civility showed marked improvements in diversity. In terms of value alignment, CVC-generated options were preferred by seven mainstream LLMs in over 70.5% of cases, surpassing both SC101 and MIC. Furthermore, CVC achieved an alignment rate exceeding 87.5% with Chinese human annotators, confirming its strong representation of Chinese cultural values.
@@
    Proposing a rule-driven method for large-scale moral dilemma generation: Leveraging CVC, we propose an automated, method to generate complex moral dilemmas based on value priorities. This approach efficiently created 404,505 dilemmas, with a subset of 10,998 evaluated across 17 LLMs, yielding 7,191 instances of divergent responses. The results demonstrate CVC’s effectiveness in generating diverse, nuanced scenarios, offering a scalable and cost-effective solution for evaluating value preferences in LLMs.
2 Related work
@@
3 Chinese Values Corpus
In this work, we present first large-scale, curated Chinese Values Corpus (CVC), designed to systematically support the alignment and evaluation of LLMs within the Chinese value system. First, based on the core socialist values, we propose a structured and hierarchical value framework (Section 3.1), which includes 12 core values and 50 derived values across national, societal, and personal dimensions. Building on this framework, we design a systematic data construction pipeline and collect value-related data from two major sources: curated international rule corpora and Chinese contexts (Section 3.2). Subsequently, we design a standardized rule-writing template and carry out large-scale rule construction, incorporating human-annotated quality control to ensure the accuracy, generalizability, and value alignment of the resulting rules (Section 3.3).
@@ 3.1 Chinese Value Framework
3.2 Data Sources of CVC
As shown in Figure 3, the CVC is built from two main data sources. The first includes curated international rule corpus, such as SC101forbes2020social and MICziems2022moral . After filtering through our value framework, basic value rules related to concepts like friendliness and integrity from these datasets can be used to expand the scale of the rule collection. The second source consists of value rules rooted in the Chinese cultural context, such as those reflecting democracy, civility, and rule of law, which ensure the dataset’s domestic relevance.
Figure 3: The overall process of constructing CVC. The blue box represents the data filtering and selection process of rules from SC101forbes2020social and MICziems2022moral , while the orange box depicts the process of constructing rules based on the Chinese cultural context. A detailed description of the construction process can be found in Appendix A.
@@ 3.2 Data Sources of CVC
The Chinese value rules are drawn from three sources: (1) academic datasets such as FLAMEShuang2023flames and the Chinese Moral Sentence Datasetpeng2021morality ; (2) existing crawled corpora such as Zhihu-KOL and People’s Daily; and (3) additional data collected via web crawling from platforms like Zhihu, Tieba, People’s Daily, and Xuexi Qiangguo. Once basic scenarios are collected, Qwen2.5-72B is again employed using a filtering template (see Appendix A.3) to remove scenarios clearly unrelated to values. After three rounds of filtering and TF-IDF-based deduplicationsalton1975vector , we obtain a total of 232,572 basic scenarios. Detailed statistics are presented in Table 6 in Appendix A.
3.3 Construction of the CVC
@@
Based on the Chinese value framework, we assign each rule a value attribute ranging from abstract to concrete. These attributes serve as essential metadata for downstream annotation tasks and scenario generation for evaluation. They help organize the CVC in a more structured and systematic manner. We employ a LLM to perform the value attribute classification. Details of this classification process are provided in Appendix A.5.
@@
Currently, mainstream LLMs exhibit limited alignment with Chinese values and norms liu2023alignbench ; khan2022ethics . While LLMs can assist in data processing, their outputs often require careful human review and adjustment to ensure full compatibility with China’s value framework. To ensure the quality and manageability of data, a representative subset of 36,000 rules was selected from the original CVC for manual annotation. Regardless of whether the rules were associated with a single derived value or multiple derived values, they were randomly sampled at a 1:6 ratio to maintain balanced coverage.
@@
Forty annotators—including experts in AI and philosophy—were recruited and trained to ensure a consistent understanding of the annotation task. To reduce individual bias, each rule was independently annotated by two annotators; in cases of disagreement, a third annotator was consulted. The final decision was made based on majority vote. The annotation tasks included identifying whether a rule was unrelated to values, inconsistent with the Chinese value system, or semantically incomplete, and rewriting the rule when necessary.
@@
Based on the manual annotations, 1,980 rules were identified as unrelated to values, 204 were found to violate the Chinese value system, and 1,426 were rewritten, yielding a total of 34,020 high-quality rules. Subsequently, five rules were randomly sampled from each derived value and further annotated by LLMs. Following both human and LLM annotations, 259,111 rules were retained, including 39,839 rewritten rules. To enhance diversity, we applied TF-IDFsalton1975vector to remove rules with a similarity score above 0.9, resulting in a final set of 257,609 high-quality and diverse rules. Detailed procedures are provided in Appendix A.6.
Statistical Overview of the CVC
@@
Figure 10 in Appendix A shows the distribution of rules across the three value levels in the CVC. Although personal-level rules dominate—mainly due to the individual-focused nature of SC101forbes2020social and MICziems2022moral —all core values across national, societal, and personal domains are fully covered. While national and societal-level rules are relatively limited, they can be supplemented through LLM-based generation (which falls beyond the scope of this study). The current dataset sufficiently supports value-alignment research in the Chinese context.
4 Empirical Validation of CVC for LLM Value Alignment
4.1 Task I: CVC-Guided Scenario Generation
@@
To accurately assess the values of LLMs, evaluation scenarios must have clear value direction and semantic boundaries. However, existing methods often rely on models generating scenarios freely, leading to issues like vague themes and value confusion. This section demonstrates the role of CVC in guiding scenario generation. By comparing the relevance and diversity of scenarios generated with and without CVC guidance, we evaluate CVC’s effectiveness in improving generation quality and clarifying value expression.
@@
To evaluate the thematic relevance of the generated scenes, we employed t-SNE van2008visualizing for dimensionality reduction and visualization, as shown in Figure 13 within Appendix B. We found that rule-guided scenarios exhibit clearer semantic boundaries and broader distribution, with minimal overlap across values. In contrast, unguided scenarios show significant intermixing. These results indicate that rule guidance improves scenario alignment with target values, enhancing their thematic relevance for LLM alignment evaluation.
@@
Value Without Rule With CVC Diff Value Without Rule With CVC Diff
@@
4.2 Task II: Chinese Value Alignment of CVC vs. Existing Benchmarks
Existing value rule benchmarks are primarily developed within Western cultural contexts. While generally applicable, they often lack the capacity to capture culturally grounded value expressions and context-specific moral reasoning in Chinese society. To evaluate the Chinese value alignment and generalizability of CVC, we select six sensitive themes and construct evaluation tasks in which CVC, SC101forbes2020social , and MICziems2022moral are used to generate options. By conducting consistency assessments across multiple mainstream language models, we examine the cultural relevance and normative coherence of different value systems, highlighting the strengths of CVC in reflecting Chinese values and supporting broader value alignment.
@@
From the themes displayed in Figure 1 we selected six representative issues—surrogacy, drugs, prejudice, firearms, politics, and suicide—and constructed corresponding value rule pairs from the CVC, SC101, and MIC datasets. Comprehensive information on the selected value-rule pairs is presented in Table 10, Appendix C. For each value pair, five scenarios were generated, each accompanied by a set of response options. The option set includes one option aligned with the CVC rule, one aligned with either SC101forbes2020social or MICziems2022moral , and one distractor. Using Qwen2.5-72B, we generated a small-scale test set containing 170 scenarios. The distribution of scenarios across different themes is also detailed in Table 12, Appendix C.
@@
Seven representative LLMsyang2024qwen2 ; hurst2024gpt ; liu2024deepseek ; grattafiori2024llama ; team2024gemini were evaluated in this study. Detailed descriptions of these models are provided in Appendix F. Each model was tested five times per theme, and the average selection rate of the CVC-aligned option was used as the model’s consistency score with CVC.
@@
The test set was independently annotated by five human evaluators, all of whom have a Chinese cultural background and were not involved in the construction of CVC. This ensures the objectivity and independence of the evaluation process.
@@
Since one of the options in each generated scenario is derived from the CVC rule, we use the selection rate of the CVC option as the primary metric to assess the alignment between models (and humans) and CVC.

### v5

# C-VARC: A Large-Scale Chinese Value Rule Corpus for Value Alignment of Large Language Models

@@ Abstract
Ensuring that Large Language Models (LLMs) align with mainstream human values and ethical norms is crucial for the safe and sustainable development of AI. Current value evaluation and alignment are constrained by Western cultural bias and incomplete domestic frameworks reliant on non-native rules; furthermore, the lack of scalable, rule-driven scenario generation methods makes evaluations costly and inadequate across diverse cultural contexts. To address these challenges, we propose a hierarchical value framework grounded in core Chinese values, encompassing three main dimensions, 12 core values, and 50 derived values. Based on this framework, we construct a large-scale Chinese Value Rule Corpus (C-VARC) containing over 250,000 value rules enhanced and expanded through human annotation. Experimental results demonstrate that scenarios guided by C-VARC exhibit clearer value boundaries and greater content diversity compared to those produced through direct generation. In the evaluation across six sensitive themes (e.g., surrogacy, suicide), seven mainstream LLMs preferred C-VARC generated options in over 70.5% of cases, while five Chinese human annotators showed an 87.5% alignment with C-VARC, confirming its universality, cultural relevance, and strong alignment with Chinese values. Additionally, we construct 400,000 rule-based moral dilemma scenarios that objectively capture nuanced distinctions in conflicting value prioritization across 17 LLMs. Our work establishes a culturally-adaptive benchmarking framework for comprehensive value evaluation and alignment, representing Chinese characteristics.
1 Background & Summary
In recent years, Large Language Models (LLMs) have been widely deployed across various domains, exerting growing influence on human decision-making and behaviortessler2024ai; hu2025generative. However, their outputs still pose significant risks, including harmful biashu2025generative, hallucinationsun2024ai; zhou2024larger, and factual inconsistencylin2022truthfulqa. These risks highlight the urgent need for effective evaluation frameworks that can assess and guide the ethical behavior of LLMs, ensuring that their responses remain aligned with prevailing societal values and moral expectations.(shen2023large; taddeo2018ai).
@@ 1 Background & Summary
To address these issues, in this paper, we propose a hierarchical value framework based on core Chinese values. With the assistance of LLMs and manual verification, we constructed a large-scale, carefully curated and standardized Chinese value corpus containing over 250,000 rules. The proposed C-VARC can effectively guide the generation of value assessment scenarios, demonstrating significant advantages in thematic relevance, value boundaries, content diversity, and semantic clarity. Furthermore, with high-quality human annotation ensuring alignment with Chinese cultural context, C-VARC provides a benchmark that both represents local values and connects with the universal ethical principles recognized by existing LLMs. Compared to other benchmarks, it provides more localized and precise standards for value alignment. Finally, C-VARC enables automated construction of complex, large-scale moral dilemma scenarios that systematically avoid biases while capturing the trade-offs between conflicting values.
@@
    The first large-scale, curated Chinese Value Rule Corpus (C-VARC): Based on the core socialist values, we develop a localized value framework covering national, societal, and personal dimensions, with 12 core values and 50 derived values. Building upon this framework, we build the first large-scale Chinese value corpus, comprising over 250,000 high-quality, manually annotated and augmented normative rules.
@@
    Systematic validation of C-VARC’s generative advantages and value alignment: We validated the effectiveness of C-VARC across all 12 core value dimensions, where C-VARC guided scenarios demonstrated clearer semantic boundaries and better category separation.Notably, C-VARC guided scenarios in categories such as rule of law and civility showed marked improvements in diversity. In terms of value alignment, C-VARC generated options were preferred by seven mainstream LLMs in over 70.5% of cases, surpassing both SC101 and MIC. Furthermore, C-VARC achieved an alignment rate exceeding 87.5% with Chinese human annotators, confirming its strong representation of Chinese cultural values.
@@
    Proposing a rule-driven method for large-scale moral dilemma generation: Leveraging C-VARC, we propose an automated, method to generate complex moral dilemmas based on value priorities. This approach efficiently created 404,505 dilemmas, with a subset of 10,998 evaluated across 17 LLMs, yielding 7,191 instances of divergent responses. The results demonstrate C-VARC’s effectiveness in generating diverse, nuanced scenarios, offering a scalable and cost-effective solution for evaluating value preferences in LLMs.
Related Work
@@
2 Methods
In this work, we present first large-scale, curated Chinese Value Rule Corpus (C-VARC), designed to systematically support the alignment and evaluation of LLMs within the Chinese value system. First, based on the core socialist values, we propose a structured and hierarchical value framework (Section 2.1), which includes 12 core values and 50 derived values across national, societal, and personal dimensions. Building on this framework, we design a systematic data construction pipeline and collect value-related data from two major sources: curated international rule corpora and Chinese contexts (Section 2.2). Subsequently, we design a standardized rule-writing template and carry out large-scale rule construction, incorporating human-annotated quality control to ensure the accuracy, generalizability, and value alignment of the resulting rules (Section 2.3).
@@ 2.1 Chinese Value Framework
2.2 Data Sources of C-VARC
As shown in Figure 3, the C-VARC is built from two main data sources. The first includes curated international rule corpus, such as SC101forbes2020social and MICziems2022moral. After filtering through our value framework, basic value rules related to concepts like friendliness and integrity from these datasets can be used to expand the scale of the rule collection. The second source consists of value rules rooted in the Chinese cultural context, such as those reflecting democracy, civility, and rule of law, which ensure the dataset’s domestic relevance.
Figure 3: The overall process of constructing C-VARC. The blue box represents the data filtering and selection process of rules from SC101forbes2020social and MICziems2022moral, while the orange box depicts the process of constructing rules based on the Chinese cultural context. A detailed description of the construction process can be found in Appendix A.
@@ 2.2 Data Sources of C-VARC
The Chinese value rules are drawn from three sources: (1) academic datasets such as FLAMEShuang2023flames and the Chinese Moral Sentence Datasetpeng2021morality; (2) existing crawled corpora such as Zhihu-KOL and People’s Daily; and (3) additional data collected via web crawling from platforms like Zhihu, Tieba, People’s Daily, and Xuexi Qiangguo. Once basic scenarios are collected, Qwen2.5-72B is again employed using a filtering template (see Appendix A.3) to remove scenarios clearly unrelated to values. After three rounds of filtering and TF-IDF-based deduplicationsalton1975vector, we obtain a total of 232,572 basic scenarios. Detailed statistics are presented in Table 1.
Table 1: Distribution of sources for basic scenario data
Source Data Volume
Zhihu-KOL 4,847
People’s Daily 8,840
Flames 671
Chinese Moral Sentence Dataset 28,536
Encyclopedia QA (JSON version) 11,622
Web Crawling 178,056
2.3 Construction of the C-VARC
@@
Based on the Chinese value framework, we assign each rule a value attribute ranging from abstract to concrete. These attributes serve as essential metadata for downstream annotation tasks and scenario generation for evaluation. They help organize the C-VARC in a more structured and systematic manner. We employ a LLM to perform the value attribute classification. Details of this classification process are provided in Appendix A.5.
@@
Currently, mainstream LLMs exhibit limited alignment with Chinese values and norms liu2023alignbench; khan2022ethics. While LLMs can assist in data processing, their outputs often require careful human review and adjustment to ensure full compatibility with China’s value framework. To guarantee quality and manageability, we selected a representative subset of 36,000 rules from the original C-VARC for manual annotation, with random sampling at a 1:6 ratio across single- and multi-valued rules to maintain balanced coverage.
A total of 40 trained annotators with backgrounds in philosophy and AI participated in the annotation. Recruitment was conducted through open calls, and all participants signed informed consent forms. Prior to the task, they received in-person training sessions covering objectives, annotation templates, and tool usage. To ensure consistency, structured annotation guidelines were provided, requiring that (1) each rule must exhibit a clear value orientation aligned with at least one core or derived value; (2) the content must be constructive, free of hate speech, violence, or discrimination; (3) the rule must be semantically complete, coherent, and logically consistent; and (4) the rule must not contradict existing laws, regulations, or mainstream norms.
@@
Each rule was independently annotated by two annotators. In cases of disagreement, a third annotator served as an arbiter, and the final decision was based on majority vote. Inter-annotator agreement measured by Cohen’s $\kappa$ exceeded 0.85, indicating high reliability. The annotation tasks involved identifying whether a rule was unrelated to values, inconsistent with the Chinese value system, or semantically incomplete, with rewriting performed when necessary.
@@
From this process, 1,980 rules were marked as unrelated to values, 204 as inconsistent, and 1,426 were rewritten, yielding 34,020 high-quality rules. To further enhance scalability, five rules from each derived value were re-annotated using LLMs trained on the human-labeled subset. After combining human and LLM annotations, 259,111 rules were retained, including 39,839 rewritten ones. Finally, to reduce redundancy, we applied TF-IDFsalton1975vector to filter rules with similarity above 0.9, producing a final set of 257,609 high-quality and diverse rules. Detailed procedures are provided in Appendix A.6.
Statistical Overview of the C-VARC
@@
Figure 4 shows the distribution of rules across the three value levels in the C-VARC. Although personal-level rules dominate—mainly due to the individual-focused nature of SC101forbes2020social and MICziems2022moral—all core values across national, societal, and personal domains are fully covered. While national and societal-level rules are relatively limited, they can be supplemented through LLM-based generation (which falls beyond the scope of this study). The current dataset sufficiently supports value-alignment research in the Chinese context.
Figure 4: Distribution of data across the three value levels in the C-VARC.
3 Data Records
The dataset used in this study is available at https://huggingface.co/datasets/Beijing-AISI/C-VARC. The Chinese Value Rule Corpus (C-VARC) consists of over 250,000 value rules, structured into three main dimensions, twelve core values, and fifty derived values. The data is organized into categories based on these core values and provided in structured JSON Lines format for easy parsing and integration into various applications, including scenario generation and value alignment evaluation. Additionally, 10,000 value rules have been successfully translated into English, with further translations and data updates planned as the annotation process progresses. The dataset is hosted on the Hugging Face platform, and detailed documentation and rule-based scenario examples are included for academic use.
4 Technical Validation
4.1 Task I: C-VARC Guided Scenario Generation
@@
To accurately assess the values of LLMs, evaluation scenarios must have clear value direction and semantic boundaries. However, existing methods often rely on models generating scenarios freely, leading to issues like vague themes and value confusion. This section demonstrates the role of C-VARC in guiding scenario generation. By comparing the relevance and diversity of scenarios generated with and without C-VARC guidance, we evaluate C-VARC’s effectiveness in improving generation quality and clarifying value expression.
@@
To evaluate the thematic relevance of the generated scenes, we employed t-SNE van2008visualizing for dimensionality reduction and visualization, as shown in Figure 5. We found that rule-guided scenarios exhibit clearer semantic boundaries and broader distribution, with minimal overlap across values. In contrast, unguided scenarios show significant intermixing. These results indicate that rule guidance improves scenario alignment with target values, enhancing their thematic relevance for LLM alignment evaluation.
@@
Value w/o Rule w/ C-VARC Diff Value w/o Rule w/ C-VARC Diff
@@
4.2 Task II: Chinese Value Alignment of C-VARC vs. Existing Benchmarks
Existing value rule benchmarks are primarily developed within Western cultural contexts. While generally applicable, they often lack the capacity to capture culturally grounded value expressions and context-specific moral reasoning in Chinese society. To evaluate the Chinese value alignment and generalizability of C-VARC, we select six sensitive themes and construct evaluation tasks in which C-VARC, SC101forbes2020social, and MICziems2022moral are used to generate options. By conducting consistency assessments across multiple mainstream language models, we examine the cultural relevance and normative coherence of different value systems, highlighting the strengths of C-VARC in reflecting Chinese values and supporting broader value alignment.
@@
From the themes displayed in Figure 1 we selected six representative issues—surrogacy, drugs, prejudice, firearms, politics, and suicide—and constructed corresponding value rule pairs from the C-VARC, SC101, and MIC datasets. Comprehensive information on the selected value-rule pairs is presented in Table 13, Appendix C. For each value pair, five scenarios were generated, each accompanied by a set of response options. The option set includes one option aligned with the C-VARC rule, one aligned with either SC101forbes2020social or MICziems2022moral, and one distractor. Using Qwen2.5-72B, we generated a small-scale test set containing 170 scenarios. The distribution of scenarios across different themes is also detailed in Table 3.
Table 3: Number of evaluation scenarios per theme
Theme Number of Scenarios
Surrogacy 35
Prejudice 5
Politics 50
Firearms 35
Drugs 15
Suicide 30
@@
Seven representative LLMsyang2024qwen2; hurst2024gpt; liu2024deepseek; grattafiori2024llama; team2024gemini were evaluated in this study. Detailed descriptions of these models are provided in Appendix F. Each model was tested five times per theme, and the average selection rate of the C-VARC aligned option was used as the model’s consistency score with C-VARC.
@@
The test set was independently annotated by five human evaluators, all of whom have a Chinese cultural background and were not involved in the construction of C-VARC. This ensures the objectivity and independence of the evaluation process.
@@
Since one of the options in each generated scenario is derived from the C-VARC rule, we use the selection rate of the C-VARC option as the primary metric to assess the alignment between models (and humans) and C-VARC.